# Feasibility of Federated Learning from Client Databases with Different Brain Diseases and MRI Modalities

### v1

@@ Abstract
Segmentation models for brain lesions in MRI are commonly developed for a specific disease and trained on data with a predefined set of MRI modalities. Each such model cannot segment the disease using data with a different set of MRI modalities, nor can it segment any other type of disease. Moreover, this training paradigm does not allow a model to benefit from learning from heterogeneous databases that may contain scans and segmentation labels for different types of brain pathologies and diverse sets of MRI modalities. Is it feasible to use Federated Learning (FL) for training a single model on client databases that contain scans and labels of different brain pathologies and diverse sets of MRI modalities? We demonstrate promising results by combining appropriate, simple, and practical modifications to the model and training strategy: Designing a model with input channels that cover the whole set of modalities available across clients, training with random modality drop, and exploring the effects of feature normalization methods.  Evaluation on 7 brain MRI databases with 5 different diseases shows that such FL framework can train a single model that is shown to be very promising in segmenting all disease types seen during training. Importantly, it is able to segment these diseases in new databases that contain sets of modalities different from those in training clients.  These results demonstrate, for the first time, feasibility and effectiveness of using FL to train a single segmentation model on decentralised data with diverse brain diseases and MRI modalities, a necessary step towards leveraging heterogeneous real-world databases. Code will be made available at: https://github.com/FelixWag/FL-MultiDisease-MRI
Keywords:
multi-modal MRI brain lesions federated learning
@@ 1 Introduction
Brain MRI segmentation is crucial for disease analysis, diagnosis and treatment. Traditionally, brain MRI segmentation models are trained on data for a specific disease, acquired with a predefined protocol and set of MRI modalities [8, 9, 30, 28]. Models trained with this standard approach cannot segment other types of diseases, and expect the same modalities in all training and testing data. While a recent study [26] demonstrated the feasibility of training a segmentation model across multiple brain MRI databases with different diseases and sets of modalities, that study used a centralised database. It did not address the practical challenge related to the sensitive nature of medical data, which prevents data sharing between institutions and therefore the creation of an appropriately diverse centralised database due to legal and privacy concerns.
@@
To address this, this study aims to develop a Federated Learning (FL) framework [15] for training a single model on diverse brain lesion databases, each with a different disease and set of MRI modalities. Central questions for this study are: How to make such FL training feasible? Will it be beneficial compared to training a disease-specific model? Will it result in a model that can segment new data that have different combination of modalities than the training data?
Related Work: While previous studies investigated training with heterogeneous databases [27, 22, 16], they have not studied FL using decentralised databases, nor data with diverse sets of MRI modalities. A recent study [26] investigated heterogeneous databases with diverse sets of MRI modalities in a centralized scenario. In contrast, our study focuses on the challenging setting of decentralized FL, which has not been explored yet. Some works have focused on FL of a segmentation model for a single, specific disease when a specific set of multi-modal MRI is available in each client [11, 19, 17]. Other works studied FL for learning from data with different labels in segmentation of CT [20] or classification in MRI [6]. In summary, there are works that focus on FL with multi-modal MRI and FL for heterogeneous tasks; however, the combination of both, especially with different sets of modalities per client, has never been explored.
Learning from databases with varying sets of MRI modalities relates to research on developing multi-modal models that perform well even when a modality is missing at inference time [4, 18, 21, 5, 31, 24]. This line of work focuses on learning a model for a specific disease, from data with the same set of modalities during training. While addressing missing modalities is an important capability of our method, it is not the primary topic of this research. Instead, this study concentrates on the effectiveness and practicality of learning from diverse databases with different sets of input modalities and segmentation tasks.
@@
Contribution: This study develops a simple but practical solution for training via FL a single model that can process different sets of modalities and segment brain lesions of different types. We show this can be achieved by a standard U-net, enhanced such that input channels can receive the whole set of unique modalities available across clients and using random drop of modalities during training to prevent associations of modalities with specific pathologies. Additionally, due to the high data distribution shift across diverse databases, we present an analysis of the effects of feature normalization techniques. Experimentation on 7 brain MRI databases with 5 different pathologies and 6 MRI modalities provides the following findings:
@@
  • Segmenting new combination of modalities: This study shows that models trained with this FL paradigm can segment data from new databases with combination of modalities not seen during training.
Figure 1: Method overview: (1) Initialization of model with distinct set of modalities across clients. (2) Federated training of the model with random modality drop. (3) Generalization to unseen clients with different input MRI modality combination.
@@ 2 Method
2.1 Problem setting: Multi-Modal MRI FL for Multiple Tasks:
We consider $C$ multi-modal clients, where each client $c$ ($1\leq c\leq C$) is associated with a set of MRI-modalities $\mathcal{M}_{c}$ and a dataset $\mathcal{D}_{c}=\{(\mathcal{X}_{ci},y_{ci})\}_{i=1}^{|\mathcal{D}_{c}|}$, where $\mathcal{X}_{ci}=\{{x}_{ci}^{m}\}_{m=1}^{|\mathcal{M}_{c}|}$ denotes the multi-modal MRI image and $y_{ci}$ the corresponding segmentation mask of the $i$-th sample. Here, each $x_{ci}^{m}\in\mathbb{R}^{w\times h\times d}$, represents a 3D image. Furthermore, $\mathcal{M}=\bigcup_{i=1}^{C}M_{i}$ denotes the set that represents all unique modalities across all clients and $p:=|\mathcal{M}|$ is the total number of unique modalities across all datasets. Each client solves a specific segmentation task $\mathcal{T}_{c}$. Our goal is train, without data sharing, a single global model parameterized by $\theta$, that performs well on all tasks and clients. The global objective is to minimize the sum of the client-local loss functions:
@@ 2.1 Problem setting: Multi-Modal MRI FL for Multiple Tasks:
where $\mathcal{L}_{c}$ denotes the local loss function of client $c$. As loss functions, we use a weighted combination of Dice and pixel-wise Binary Cross-Entropy (BCE) losses $\mathcal{L}_{c}=\alpha\mathcal{L}_{Dice}+(1-\alpha)\mathcal{L}_{BCE}$, where $\alpha=0.8$ in our experiments. In the standard FL setting with $E$ communication rounds, a global server initializes a global model and communicates it to every client. In each round, each client performs $\tau$ iterations of updates on its local data and loss function $\mathcal{L}_{c}(\theta;\mathcal{D}_{c};\mathcal{T}_{c})$. Afterwards, every client sends their updated model back to the server. The server combines these updates, typically by averaging (FedAvg), into a new global model before starting a new communication round [15]. This allows decentralised training without data sharing. We want to highlight the differences from the standard FL setting: this setting varies not only in the input space $|\mathcal{M}|>1$, indicating multiple modalities, but also in the label space $|\bigcup_{i=1}^{C}T_{i}|>1$, involving segmenting different brain pathologies per client. Additionally, this setting considers different sets of modalities per client, an aspect mostly overlooked in FL.
2.2 Training a Unified Model for Multiple Segmentation Tasks
@@
Handling Different Sets of Modalities Across Clients: Since every client comes with a different combination of MRI modalities, the challenge arises to find a single model architecture capable of processing different sets of input modalities. To address this, we augment each client’s input to cover the full set of distinct MRI modalities across all clients, and fill missing modalities $m\in\mathcal{M}-\mathcal{M}_{c}$ of each client $c$ with zeros. This transforms the input space of each client to $\mathcal{X}_{c}\in\mathbb{R}^{p\times w\times h\times d}$. Consequently, our global model accepts the set of distinct MRI modalities as input and becomes a function $f(\cdot;\theta):\mathbb{R}^{p\times w\times h\times d}\to\mathbb{R}^{w\times h\times d}$, capable of processing each client’s input modality combination, ensuring compatibility across all clients.
@@
Modality Drop for Pathology Disassociation & Missing Modalities: To develop a model that generalises to modality combinations not seen during training, we introduce modality drop, where we randomly set input modalities to zero during training. Therefore, we define a probabilistic masking function $Z(\mathcal{X},\phi)$ where each input modality $x^{m}\in\mathcal{X}$ is set to zero with probability $\phi$. This prevents the model from associating specific modality combinations with specific pathologies, ensuring it can accurately segment known pathologies in unseen clients with modality combinations not seen during training. Furthermore, it also enhances the model’s robustness against missing modalities during testing, as shown in our evaluation.
@@
Feature Normalization for Heterogeneous Clients: Each client’s data contains different pathologies and modalities, which increases the difficulty of training a model due to data heterogeneity. To tackle this, we employ Batch Normalization (BN) [7], but keep client-specific BN parameters and statistics, and exclude them from the model averaging step, as proposed in [1, 12]. Our experiments show that this method proves effective in improving the performance of the training (source) databases. However, it is not clear how to apply the source-specific BN layer parameters/statistics to unseen (target) clients during inference. One way involves averaging the learned BN parameters and estimating running statistics on the target data, which has the disadvantage that target data is needed. Therefore, in our experiments, we also explore alternative feature normalization methods for this training framework with inherently high data heterogeneity. Specifically, the recently proposed Normalization Free (NF) technique [32] uses re-parameterized convolutional layers instead of explicitly normalizing statistics of feature activations. This does not require target data to estimate the target statistics, alleviating this requirement of keeping client-specific BN layers. Empirical analysis (Sec. 3) also shows that this results in methods that generalize better to unseen clients. However, it slightly decreases performance on source clients. Therefore, the choice of feature normalization method in this framework should be chosen based on the target application for the trained model (generalization to new domains or application to source domains) and whether data are expected to have been collected a-priori from the target domain for estimating BN statistics.
@@
Algorithm: To summarise, our simple yet effective algorithm is shown in Fig. 1. Initially, the central server receives a list of available MRI modalities from each client’s dataset. Afterwards, the server initializes a model architecture that accepts the set of unique MRI modalities $\mathcal{M}$ as input and the FL algorithm starts (Sec. 2.1). During training, each client’s input modalities are augmented to the whole set of unique MRI modalities and modality drop is applied.

### v2

@@ Abstract
Segmentation models for brain lesions in MRI are commonly developed for a specific disease and trained on data with a predefined set of MRI modalities. Each such model cannot segment the disease using data with a different set of MRI modalities, nor can it segment any other type of disease. Moreover, this training paradigm does not allow a model to benefit from learning from heterogeneous databases that may contain scans and segmentation labels for different types of brain pathologies and diverse sets of MRI modalities. Additionally, the sensitivity of patient data often prevents centrally aggregating data, necessitating a decentralized approach. Is it feasible to use Federated Learning (FL) to train a single model on client databases that contain scans and labels of different brain pathologies and diverse sets of MRI modalities? We demonstrate promising results by combining appropriate, simple, and practical modifications to the model and training strategy: Designing a model with input channels that cover the whole set of modalities available across clients, training with random modality drop, and exploring the effects of feature normalization methods.  Evaluation on 7 brain MRI databases with 5 different diseases shows that such FL framework can train a single model that is shown to be very promising in segmenting all disease types seen during training. Importantly, it is able to segment these diseases in new databases that contain sets of modalities different from those in training clients.  These results demonstrate, for the first time, the feasibility and effectiveness of using Federated Learning to train a single 3D segmentation model on decentralised data with diverse brain diseases and MRI modalities, a necessary step towards leveraging heterogeneous real-world databases. Code will be made available at: https://github.com/FelixWag/FL-MultiDisease-MRI
Figure 1: FedUniBrain: Traditionally, brain lesion MRI segmentation models are trained for a specific disease, using MRI modalities determined by a study-specific protocol. Each resulting model is capable of only processing a predefined set of MRI input modalities, segmenting a single disease and is only trained on a limited number of samples. In this study, we introduce the FL framework FedUniBrain, which collaboratively trains a single model across multiple institutions capable of segmenting multiple diseases and is not bound to a predefined set of MRI input modalities. Our framework allows decentralized training without any data sharing, respecting the private nature of medical data.
@@ 1 Introduction
Brain MRI segmentation is crucial for disease analysis, diagnosis, and treatment. Multi-modal MRI scans provide complementary insights essential for detecting diseases. Each modality provides complementary information for tissues and anatomies via unique contrasts, which is crucial for diagnosing different diseases. For instance, the common image modality FLAIR suppresses cerebrospinal fluid signal and is therefore useful for detecting periventricular lesions like Multiple Sclerosis. Models that can jointly process all available MRI modalities for a patient are therefore paramount for brain MRI analysis. Traditionally, such multi-modal MRI models for segmentation of brain lesions are trained on data for a specific disease, acquired with a predefined protocol and set of MRI modalities [16, 17, 46, 44]. Models trained with this standard approach cannot segment other types of diseases, nor can they process patient scans if a different set of MRI modalities has been acquired compared to the specific set in the training data. While a recent study [41] demonstrated the feasibility of training a segmentation neural network across multiple brain MRI databases with different diseases and sets of modalities, that study used a centralised database as proof-of-concept. It did not address the practical challenge that, due to the sensitive nature of medical data and privacy concerns, data sharing between institutions is not straight-forward, which prevents the creation of an appropriately diverse centralised database. Federated Learning (FL) [27] allows collaboratively training a model without sharing or centrally aggregating data, protecting sensitive patient information. It is an open question whether it is possible to effectively train a model with FL using decentralised databases, each with different types of pathologies and sets of MRI modalities, which could then enable larger scale training of such models via multi-institutional collaborations.
@@
To address all these challenges, this study presents the FL framework FedUniBrain (Federated Universal Brain MRI Segmentation) for training a single 3D segmentation model on diverse brain lesion databases, each with a different disease and set of MRI input modalities, as shown in Fig. 1. This represents a challenging federated optimization setting since there is not only high heterogeneity within the MRI data from different scans across various databases but also across different segmentation tasks, as well as diverse MRI input modalities. While a few recent FL approaches have been developed to handle heterogeneous tasks and modalities per client, none of them is applied in the medical domain to address multiple MRI modalities per client for 3D segmentation. Central questions for this study are: Is federated training of such a model feasible? Will it be beneficial compared to training a disease-specific model? Will it result in a model that can segment new data with a different combination of modalities than the training data? Can a new client join in the training with a modality and/or pathology not seen during training?
@@
Contribution: This study develops a simple but practical solution for training via FL a single model that can process different sets of modalities and segment brain lesions of different types. We show that this can be achieved by using a standard U-net, enhanced such that input channels can receive the whole set of unique modalities available across clients, and using random drop of modalities during training to prevent associations of modalities with specific pathologies. Additionally, due to the high data distribution shift across diverse databases, we present an analysis of the effects of feature normalization techniques. Experimentation on 7 brain MRI databases with 5 different pathologies and 7 MRI modalities provides the following findings:
@@
  • Segmenting new combination of modalities: This study shows that models trained with this paradigm can segment data from new databases with combination of modalities not seen during training.
  • Client with unseen modality and/or pathology joins the federation: We show that our framework is flexible and able to handle clients that join the federation with a previously unseen modality and/or pathology.
2 Related Work
Medical Image Segmentation on Heterogeneous Databases. Previous studies on medical image segmentation/lesion detection explored training with heterogeneous databases [42, 37, 28, 5]. However, these studies focus on centralised training, where data from different sources is combined into a single dataset. This overlooks practical real-world issues in the medical field, such as data sharing restrictions and the unique challenges associated with decentralized optimization. Additionally, they did not address the complexities of training with data with diverse sets of MRI modalities. A recent study [41] examined heterogeneous databases with diverse sets of MRI modalities in a centralized scenario. In contrast, our study focuses on the challenging setting of decentralized FL, which has not been explored yet.
Previous works on FL for brain MRI segmentation [20, 34, 29] focused on single-disease scenarios, where each client has access to the same set of MRI input modalities. These studies demonstrated that FL can effectively train a brain MRI segmentation without centralizing data. However, they did not address the challenges of diverse sets of MRI modalities and segmentation tasks per client.
Federated Multi-Task Learning. Traditional Multi-Task learning trains a model on multiple tasks simultaneously to extract shared knowledge and generalize across multiple tasks [7, 11, 30]. Early research on Federated Multi-Task Learning mainly addresses data heterogenity across clients within a Multi-Task learning framework. Recent approaches [6, 24, 9] in the non-medical domain focus on both data heterogenity as well as task heterogenity per client (e.g. each client specializes on a different segmentation task). However, none of these works is applied to medical imaging and applicable to multi-MRI modalities inputs. In medical imaging, there exists only a few papers that tackle Multi-Task FL. [35] studied FL for learning from data with different labels in segmentation of CT. [14] tackles the problem of multi-task classification of mental disorders from MRI scans. However, both of theses works only consider one imaging modality as input. In summary, while there are works on FL with multi-modal MRI and FL for heterogeneous tasks; the combination of both, especially with different sets of modalities per client, has not been explored.
Missing Modalities. Learning from databases with varying sets of MRI modalities relates to research of developing multi-modal models that perform well even when a modality is missing at inference time [12, 33, 36, 13, 47, 39]. Missing modalities in clinical practice are often caused by acquisition protocols, scanning costs or scanner availability. This line of work focuses on learning a model for a specific disease, from data with the same set of modalities during training. Furthermore, these methods are trained in centralised settings. FL for missing modalities is an emerging research direction. Most existing studies focus on text and image modalities [31, 4], non-imaging data [8], or require unimodal or complete multimodal data [43]. Notably, none of these approaches have been applied to different MRI modalities to segment multiple tasks. While addressing missing modalities is an important capability of our method, it is not the primary topic of this research. Instead, this study concentrates on the effectiveness and practicality of learning from diverse databases with different sets of input modalities and segmentation tasks.
Comparison to Foundation Models. Our study has some similarities with the area of foundation models that are able to perform well and generalize to a variety of tasks, like SAM [18] and especially for medical imaging MedSAM [25]. However, there is a clear distinction to our work. Those models are not designed to handle 3D input data as well receive multi modality MRI input. Additionally, our method is not an interactive method that requires any user input prompts. Therefore, this line of research is not directly applicable to our problem setting.
Figure 2: Method overview. (1) Initialization of model with distinct set of modalities across clients. (2) Federated training of the model with random modality drop. (3) Generalization to unseen clients with different input MRI modality combination.
@@ 3 Method
3.1 Problem setting: FL with Multi-Modal MRI and Multiple Tasks
We consider $C$ clients, where each client $c$ ($1\!\leq\!c\!\leq\!C$) is associated with a multi-modal MRI dataset $\mathcal{D}_{c}=\{(\mathcal{X}_{ci},y_{ci})\}_{i=1}^{|\mathcal{D}_{c}|}$ with set of MRI-modalities $\mathcal{M}_{c}$, where $\mathcal{X}_{ci}=\{{x}_{ci}^{m}\}_{m=1}^{|\mathcal{M}_{c}|}$ denotes the multi-modal MRI image and $y_{ci}$ the corresponding segmentation mask of the $i$-th sample. Here, each $x_{ci}^{m}\in\mathbb{R}^{w\times h\times d}$, represents a 3D image, where $w,h$ and $d$ denote the width, height, and depth respectively. Furthermore, $\mathcal{M}=\bigcup_{i=1}^{C}M_{i}$ denotes the set that represents all unique modalities across all clients and $|\mathcal{M}|$ is the total number of unique modalities across all datasets. Each client solves a specific segmentation task $\mathcal{T}_{c}$. Our goal is to train, without data sharing, a single global model parameterized by $\theta$, that performs well on all tasks and clients. The global objective is to minimize the average of the client-local loss functions:
@@ 3.1 Problem setting: FL with Multi-Modal MRI and Multiple Tasks
where $\mathcal{L}_{c}$ denotes the local loss function of client $c$. As loss functions, we use a weighted combination of Dice and pixel-wise Binary Cross-Entropy (BCE) losses:
$\mathcal{L}_{c}=\alpha\mathcal{L}_{Dice}+(1-\alpha)\mathcal{L}_{BCE}$ (2)
@@
In the standard FL setting with $E$ communication rounds, a global server initializes a global model and communicates it to every client. In each round, each client performs $\tau$ iterations of updates on its local data and loss function $\mathcal{L}_{c}(\theta;\mathcal{D}_{c};\mathcal{T}_{c})$. Afterwards, every client sends their updated model back to the server. The server combines these updates, typically by averaging (FedAvg), into a new global model before starting a new communication round [27]. This allows decentralised training without data sharing.
Difference to other FL settings: We highlight that, in comparison to the standard FL setting, the setting studied herein varies not only with respect to input space with $|\mathcal{M}|\!>\!1$, considering multiple modalities, but also the label space $|\bigcup_{i=1}^{C}\mathcal{T}_{i}|\!>\!1$, segmenting different brain pathologies per client. Moreover, this setting considers different sets of modalities per client, an aspect mostly overlooked in FL. Comparison of different FL settings is shown in Tab. 3.
Table 1: Comparison between FL settings: Previous FL settings are not designed to simultaneously handle different tasks per client, multiple MRI input modalities, missing modalities, and varying sets of modality combinations across client. FedUniBrain is designed to handle all these aspects simultaneously.
Setting Multi Task Multi MRI Mod. Missing Mod. Unseen Mod. Comb.
Standard FL
Multi-Task FL
Multi-MRI Mod. FL
FedUniBrain
3.2 Unified Model for Multiple Segmentation Tasks
@@
Handling Different Sets of Modalities Across Clients: Since every client comes with a different combination of MRI modalities, the first challenge arising is to define a single neural network architecture that can process different sets of input modalities. To address this, we augment each client’s input images to cover the full set of $|\mathcal{M}|$ distinct MRI modalities across all clients. For each client $c$, any missing modalities in the set $\mathcal{M}-\mathcal{M}_{c}$ are filled with zeroes. Therefore, the input space for each client $c$ is extended from $\mathcal{X}_{c}\in\mathbb{R}^{|\mathcal{M}_{c}|\times w\times h\times d}$ (only client-specific modalities) to $\mathcal{X}_{c}\in\mathbb{R}^{|\mathcal{M}|\times w\times h\times d}$, where $|\mathcal{M}|$ is the total number of modalities across all clients. Consequently, our model accepts the set of $|\mathcal{M}|$ distinct MRI modalities as input and can be represented as a function $f(\cdot;\theta):\mathbb{R}^{|\mathcal{M}|\times w\times h\times d}\to\mathbb{R}^{w\times h\times d}$, capable of processing each client’s input modality combination and outputting a 3D segmentation mask, ensuring compatibility across all clients.
@@
Modality Drop for Pathology Disassociation & Missing Modalities: To develop a model that generalises to modality combinations not seen during training, we employ modality drop, where we randomly set input modalities to zero during training. Therefore, we define a probabilistic masking function $Z(\mathcal{X},\phi)$ where each input modality $x^{m}\in\mathcal{X}$ is set to zero with probability $\phi$. We implement the following approach: During training, for each sample, we consider the set of initially available input modalities $\mathcal{M}_{c}$ from the corresponding database. We then sample a random integer $r\in[1,|\mathcal{M}|]$, ensuring that at least one modality remains. Next, we randomly select $r$ number of modalities from $\mathcal{M}_{c}$ to keep, setting the remaining modalities blank. This prevents the model from associating specific modality combinations with specific pathologies, ensuring it can accurately segment known pathologies in unseen clients with modality combinations not seen during training. Furthermore, it also enhances the model’s robustness against missing modalities during testing, as shown in our evaluation.
(a) Segmentation results on source clients.
(b) Generalization to unseen sets of modalities.
Figure 3: Example segmentation results for databases with different pathologies. (a) Segmentation outputs for five different databases, each having a different pathology to segment and different modality combination. Shown are 2D slices of the 3D input, the ground truth, predictions from individually trained U-Nets (on each database separately), and our FedUniBrain (trained on ATLAS, BRATS, MSSEG, TBI and WMH) model predictions. For each database we show a different modality. (b) Segmentation outputs from a FedUniBrain model trained on ATLAS, BRATS, MSSEG, TBI, and WMH, tested on excluded ISLES and Tumor2 databases with unseen modality combinations. Also, results from disease-specific models are shown: one trained on the BRATS (Tumor) and evaluated on Tumor2, and another trained on ATLAS and evaluated on ISLES for stroke lesions.
Feature Normalization for Heterogeneous Clients: Batch Normalization (BN) [15] is an indispensable feature normalization technique that showed various benefits in training deep neural networks [15, 32]. BN normalizes the features for each mini-batch as follows:
$\text{BN}(\mathbf{x})=\gamma\frac{\mathbf{x}-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\beta,$ (3)
where $\mu$ and $\sigma$ represent the mean and variance computed over the mini-batch, while $\gamma$ and $\beta$ are learnable parameters. Term $\epsilon$ is used for numerical stability. During inference $\mu$ and $\sigma$ are replaced with running averages $\hat{\mu}$ and $\hat{\sigma}$ computed during training to avoid reliance on mini-batches. Despite its success, it has been shown that BN does not perform well in multi-domain training as these statistics are not representative for individual domains [22]. In our setting, each client’s data comes from a different domain, with different pathologies, modalities and scanners, which increases the difficulty of training a model due to high data heterogeneity. This introduces a challenge when using BN for our setting.
@@
To tackle this, we keep client-specific BN parameters and statistics ($\gamma_{c}$, $\beta_{c}$ and $\hat{\mu}_{c}$, $\hat{\sigma}_{c}$) for each client $c$, and exclude them from FL’s model averaging step, as proposed in [1, 21]. Our experiments show that this method improves performance on the training (source) databases. However, it is unclear how to apply the source-specific BN layer parameters/statistics to unseen (target) clients during inference. One way involves using averages of learned BN parameters, $\bar{\gamma}\!=\!\frac{1}{C}\sum_{c=1}^{C}{\gamma_{c}}$ and $\bar{\beta}\!=\!\frac{1}{C}\sum_{c=1}^{C}{\beta_{c}}$, and estimating target-specific running average statistics $\hat{\mu}_{t}$ and $\hat{\sigma}_{t}$ with a forward pass over the target database. The latter poses the disadvantage that a pre-collected target database is needed prior to inference. Therefore, in our experiments, we also explore alternative feature normalization methods for this training framework with inherently high data heterogeneity. Specifically, we investigate the recently proposed Normalization Free (NF) technique [48] that uses re-parameterized convolutional layers instead of explicitly normalizing statistics of feature activations, Group Norm (GN) [40], and Instance Norm (IN) [38]. These approaches do not require target data to estimate the target statistics, alleviating this requirement of keeping client-specific BN layers. Empirical analysis (Sec. 4) also shows that this results in methods that generalize better to unseen clients. However, it slightly decreases performance on source clients. Therefore, the choice of feature normalization method in this framework should be chosen based on the target application for the trained model (generalization to new domains or personalization to source domains) and whether data are expected to have been collected a-priori from the target domain for estimating BN statistics.
Client with unseen modality joins: In a realistic FL scenario, a new client may join a federation at a later training stage, for example, by contributing data to further improve a pre-trained model. If their data includes a new modality that was not in the originally available set $\mathcal{M}$, can we learn the new modality and data? To handle this scenario, we alter the model as follows. Assume a network pre-trained before the new client joins, and let the weights of its first convolutional layer be $\mathbf{W}_{1}\!\in\!\mathbb{R}^{o\times|\mathcal{M}|\times K_{h}\times K_{w}\times K_{d}}$, where $o$ is the number of output filters, $K_{h},K_{w},K_{d}$, the height, width, and depth of the kernel, respectively. To add the new modality, we extend dimensionality of $\mathbf{W_{1}}$ to get ${\mathbf{W^{\prime}_{1}}\in\mathbb{R}^{o\times(|\mathcal{M}|+1)\times K_{h}\times K_{w}\times K_{d}}}$. We then initialize the weights of the existing channels with the pre-trained weights: ${\mathbf{W^{\prime}_{1}}[:,i,:,:,:]=\mathbf{W_{1}}[:,i,:,:,:],\forall\,i\in[1,|\mathcal{M}|].}$ For the new modality, we randomly select one of the pre-existing modalities, sampling $k\in[1,|\mathcal{M}|]$, and use its pre-trained weights to initialize the weights associated with the newly added modality: ${\mathbf{W^{\prime}_{1}}[:,|\mathcal{M}|+1,:,:,:]=\mathbf{W_{1}}[:,k,:,:,:]}$. All network parameters, including $\mathbf{W^{\prime}_{1}}$, are then trained as normal with data from all clients.
Table 2: Separate models are trained and evaluated on each database (single center), with only BN results shown as they yield the best results. In a federated setting, FedUniBrain is trained jointly on ATLAS, BRATS, MSSEG, TBI, and WMH with various feature normalization methods. Results with/out modality drop are also reported. Results (Dice score %) show the feasibility of federated training a unified model across multiple diverse brain databases, with performance comparable to or slightly outperforming disease-specific models.
Method Mod. Drop Norm ATLAS BRATS MSSEG TBI WMH Average
Single Center BN 52.8 91.9 68.3 56.1 71.5 68.1
FedUniBrain (client spec. BN params) BN 50.9 91.8 70.2 55.8 74.8 68.7
Single Center BN 52.8 91.7 66.7 53.9 69.1 66.8
FedUniBrain (avg. BN params) BN 43.1 89.4 63.2 51.4 67.3 62.9
FedUniBrain IN 49.0 90.7 66.6 52.9 71.0 66.0
FedUniBrain GN 48.1 91.0 67.3 53.7 70.4 66.1
FedUniBrain NF 51.2 91.1 64.8 54.4 72.3 66.8
FedUniBrain (client spec. BN params) BN 54.5 91.8 69.1 56.2 73.7 69.1
@@
Algorithm: To summarise, our simple yet effective algorithm is shown in Fig. 2. Initially, the central server receives a list of available MRI modalities from each client’s dataset. Afterwards, the server initializes a model architecture that accepts the set of unique MRI modalities $\mathcal{M}$ as input, and the FL algorithm starts (Sec. 3.1). During training, each client’s input modalities are extended to the whole set of unique MRI modalities, with each missing modality filled with zeroes, and modality drop is applied.